An object is acted upon by two forces 20N each .The angle between the forces is 120 degrees.The resultant force on the object has the magnitude of what


1
Expert's answer
2021-03-08T08:23:37-0500

"F_{res}=\sqrt{F_1^2+F_2^2+2F_1F_2cos\alpha},"

"F_1=F_2=F=20~N,"

"F_{res}=\sqrt{F^2+F^2-F^2}=F=20~N."